OpenStudy (anonymous):

how would i find the acerage rate of change of h(x)=-x^2+2x+2 from x=3 to x+8

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

or \[\frac{ h(x2)-h(x1) }{ x2-x1 }\]

OpenStudy (anonymous):

how would u do the bottom one

OpenStudy (anonymous):

Is that x+8 or x=8

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

OK so we make 3 = x1 and 8 =x2 so that takes care of the denominator in our equation. Do you see this?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

OK taking x1=3 we need to determine h(x1). o do this simply take the value for x(which is 3) and plug it into the equation. The result will be h(x)1. So what did you get?